OpenStudy (anonymous):

The equation of circle O is (x - 2) 2 + (y - 4) 2 = 36. Determine which of these points are on the circumference of the circle. There may be more than one correct answer. (2, 4) (8, 4) (-2, 2) (2,-2) (-4, 4) (-2,-2) (6, 6)

OpenStudy (anonymous):

?

OpenStudy (mathmate):

Well, are you able to tell me if (2,4) is on the circumference of the circle?

OpenStudy (anonymous):

no! i dont understand

OpenStudy (anonymous):

i only have like 10 min left

OpenStudy (mathmate):

(2,4) means x=2, y=4. ALL points on the circumference of the circle satisfy the relation \[ (x - 2)^2 + (y - 4)^2 = 36\] So if the point (2,4) is on the circumference, then substituting 2 and 4 for x and y respectively will result in equality. Let's try: \[ (2 - 2)^2 + (4 - 4)^2 = 0^2+0^2 \ne 36\] so (2,4) is not on the circumference of the circle. Does it make any sense to you?

OpenStudy (anonymous):

ooo yes !!!!!!! thank u

OpenStudy (mathmate):

You're welcome ! :) You'll need to check every one of the points because there are multiple solutions.

OpenStudy (anonymous):

got it!!!!!!! :)